# Multispectral State-Space Feature Fusion: Bridging Shared and Cross-Parametric Interactions for Object Detection

### v1

@@ Abstract
Modern multispectral feature fusion for object detection faces two critical limitations: (1) Excessive preference for local complementary features over cross-modal shared semantics adversely affects generalization performance; and (2) The trade-off between the receptive field size and computational complexity present critical bottlenecks for scalable feature modeling. Addressing these issues, a novel Multispectral State-Space Feature Fusion framework, dubbed MS2Fusion, is proposed based on the state space model (SSM), achieving efficient and effective fusion through a dual-path parametric interaction mechanism. More specifically, the first cross-parameter interaction branch inherits the advantage of cross-attention in mining complementary information with cross-modal hidden state decoding in SSM. The second shared-parameter branch explores cross-modal alignment with joint embedding to obtain cross-modal similar semantic features and structures through parameter sharing in SSM.
Finally, these two paths are jointly optimized with SSM for fusing multispectral features in a unified framework, allowing our MS2Fusion to enjoy both functional complementarity and shared semantic space. Benefiting from the design of the dual-branch SSM, our approach simultaneously inherits the computational efficiency and the global receptive field, significantly improving the performance of multispectral object detection. In our extensive experiments on mainstream benchmarks including FLIR, M3FD and LLVIP, our MS2Fusion significantly outperforms other state-of-the-art multispectral object detection methods, evidencing its superiority. Moreover, MS2Fusion is general and applicable to other multispectral perception tasks. We show that, even without specific design, MS2Fusion achieves state-of-the-art results on RGB-T semantic segmentation and RGB-T salient object detection, showing its generality. The source code will be available at https://github.com/61s61min/MS2Fusion.git.
@@ 1 Introduction
As demonstrated in Figure 3, our ERF analysis reveals that MS2Fusion achieves superior spatial coverage compared to both CNN and Transformer, successfully integrating local details with global context. This capability addresses the fundamental constraint of standard CNN in modeling long-range dependencies while maintaining computational efficiency.
Table 1 also provides our component replacement experiments on the MS2Fusion module in the multispectral object detection framework (Figure 5), which yield three key findings: (1) Compared to Transformer, MS2Fusion achieves a 66.6% reduction in FLOPs and 70.4% fewer parameters while delivering better detection accuracy (+8.3 mAP50 points); (2) When benchmarked against CNN baseline, it maintains 26.0% lower computational costs and 18.4% parameter reduction, while achieving significant improvements of +9.0 mAP50 points; (3) The MS2Fusion consistently outperforms both baselines across all complexity-accuracy trade-off metrics. These systematic experiments provide conclusive evidence that MS2Fusion successfully breaks the traditional efficiency-accuracy trade-off, establishing new state-of-the-art performance in multispectral object detection.
Table 1: Comparison of different multispectral feature fusion methods
GFLOPs Params (M) mAP@0.5 mAP@0.75 mAP
CNN 190.3 159.7 74.3 23.8 32.5
Transformer 421.9 440.6 75.0 24.1 33.4
MS2Fusion 140.8 130.3 83.3 33.0 40.3
@@
    The proposed MS2Fusion establishes a dual-modal collaborative learning mechanism within state space models. The approach achieves implicit feature complementarity through CP-SSM and enhances shared features via SP-SSM, effectively addressing the issue of excessive complementary feature preference.
@@
    The MS2Fusion achieves co-optimization of receptive field coverage and computational efficiency. While maintaining computational complexity comparable to CNN, its ERF surpasses that of Transformer, breaking the inherent limitations of existing architectures.
@@
    The MS2Fusion demonstrates remarkable adaptability to diverse input modalities and functions as a plug-and-play module. Its compatibility with various backbone networks and outstanding performance across different downstream tasks substantiates both its versatility and effectiveness.
@@ 3.1 State Space Model (SSM)
SSM represents a class of architectures for sequence modeling rooted in linear time-invariant systems from cybernetics. They can be seen as an integration of recurrent neural networks (RNNs) and CNN. SSM offers several advantages, including linear-time inference, parallelized training capability, and robust performance in tasks requiring long-context dependencies.
In SSM, the modeling process involves transforming a one-dimensional input sequence $x(t)$ into an intermediate state $h(t)$ via a state equation. This intermediate state then generates an output sequence $y(t)$ through an output equation. Typically, the SSM is formulated as a linear ordinary differential equation, formulated in Equation 1:
@@
where $h(t)$ denotes the state vector, $x(t)$ represents the input vector, A is the state matrix, used to update the hidden state; B is the input matrix, describing the external input; C is the output matrix, mapping the state to the output; and D is the direct transmission matrix, representing the direct coupling from input to output.
The current SSM is suitable for continuous time scenarios and, in order to apply it in deep learning scenarios, it needs to be discretized. Therefore, a fixed time interval $\Delta$ is used to sample the input sequence, and Equation 1 can be discretized into Equation 2 by using the zero-order hold principle (indicated by a horizontal bar above the corresponding variable):
@@
[…]
After discretization, the state-space model can be subjected to convolution operations via the convolution kernel $\overline{\textbf{K}}$.
${\overline{K}}$ $=\left({\textbf{C}{\overline{\textbf{B}}},\textbf{C}{\overline{\textbf{A}}\overline{\textbf{B}}},\ldots,\textbf{C}{{\overline{\textbf{A}}}}^{L-1}{\overline{\textbf{B}}}}\right)$ (3)
$y$ $=x\cdot{\overline{\textbf{K}}}+\textbf{D}\cdot x$
where $x$ represents all input elements and $x$ denotes the input sequences. Thus, Mamba can use convolutional operations for efficient parallel computation during training, and switch to an RNN-like recurrent mode for fast autoregressive inference during inference.
Although the SSM is highly effective for modeling discrete sequences, its linear time-invariant nature imposes limitations: the model parameters remain fixed regardless of the input, potentially hindering the ability to selectively focus on relevant information while disregarding irrelevant data. To address this constraint, SSM introduces dynamic parameter adjustments (denoted as $\overline{\textbf{B}},\overline{\textbf{C}},\Delta$ in Equation 2) based on current inputs. This adaptive mechanism enables Mamba to contextualize inputs dynamically, effectively filtering out irrelevant information and emphasizing pertinent inputs. As a result, Mamba can efficiently model complex interactions inherent in long sequences, enhancing its capability to handle diverse and challenging data environments.
Figure 4: Details of SSM. (For a $L\times d$ dimensional input {$x_{1},x_{2},...,x_{L}$}.)
As shown in Figure 4, the SSM processes an input sequence $\{x_{1},x_{2},...,x_{L}\}\in\mathbb{R}^{L\times d}$ step-by-step through a structured transformation. The input first passes through a linear layer, followed by recursive state updates governed by matrices A, B, C, and D. Here, A drives the hidden state transition, B maps the input to the state space, C projects the hidden state to the output, and D optionally incorporates a direct input-to-output connection. This recurrent computation generates intermediate outputs $\{y_{1},y_{2},...,y_{L}\}\in\mathbb{R}^{L\times d}$, where each $y_{i}\in\mathbb{R}^{1\times d}$ is derived sequentially. The full process is formalized in Equation 4.
$y_{i}^{j}$ $=\left(\overline{\textbf{C}}\cdot\left(\overline{\textbf{A}}\cdot h_{i}^{j}+\overline{\textbf{B}}\cdot x_{i}^{j}\right)+\textbf{D}\cdot x_{i}^{j}\right)$ (4)
$y_{i}$ $=\left[y_{i}^{1},y_{i}^{2},\ldots,y_{i}^{d}\right]\,,(i=1,2,\dots,L;j=1,2,\dots,d)$
where $h_{i}^{j}$ denotes the ith intermediate state in the jth dimension and $y_{i}^{j}$ denotes the ith output in the jth dimension.
Figure 5: Overview of the model architecture. It consists of three main stages: (1) feature extraction with two backbone networks; (2) cross-modal feature fusion of $P3$, $P4$ and $P5$ with MS2Fusion module; (3) the detection results are generated through the Neck and Head layers. In our experiments, two distinct detection heads ( CoDetr and YOLOv5) are evaluated independently.
@@ 3.2 The Proposed Model Architecture
As shown in Figure 5, the proposed framework is structured into three main stages. Initially, the backbone network extracts features independently from RGB and thermal images. Subsequently, cross-modal MS2Fusion feature fusion integrates features from different stages. Finally, object localization and regression are performed using the detection head to derive the final detection outcomes. During the feature fusion stage, we selectively merge features from three distinct levels. P3 layer captures detailed surface information, while P5 encapsulates higher-level semantic feature. By separately fusing shallow and deep features, our approach effectively concentrates on detailed information critical for cross-modal fusion, which is formulated in Equation 5:
@@
Figure 6: The overview of the MS2Fusion module. The MS2Fusion module employs a dual-branch architecture to process the features of two modalities, $F_{V}$ and $F_{T}$. CP-SSM fuses cross-modal features while preserving modality-specific details, while SP-SSM extracts shared features. These shared features are then enhanced via FF-SSM and fused with original modality features. Finally, FF-SSM performs cross-modal fusion, outputting the fused feature $F_{fused}$.
@@ 3.3 Multispectral State Space Feature Fusion (MS2Fusion)
As shown in Figure 6, the MS2Fusion module consists of three core components: CP-SSM, SP-SSM, and FF-SSM. Firstly, the CP-SSM module achieves global feature integration of RGB and thermal modalities through a dynamic parameter interaction mechanism, enhancing cross-modal contextual awareness while preserving modality-specific characteristics. This module innovatively employs an implicit parameter crossover strategy to effectively mine and reinforce complementary features between the two modalities. Secondly, the SP-SSM module systematically extracts and enhances the common representations of both modalities by constructing a shared feature space, significantly improving the quality of unimodal features. Finally, to more effectively achieve cross-modal feature fusion, we adopt the following hierarchical processing strategy: First, we construct bidirectional feature interaction channels through two parallel FF-SSM modules, leveraging shared features to enhance the feature representation of each branch. Subsequently, we introduce a third FF-SSM module specifically dedicated to cross-modal feature fusion. This phased processing approach enables the adaptive integration of complementary and shared features across modalities, thereby significantly improving fusion performance.
The innovation of MS2Fusion lies in its systematic utilization of the dual characteristics of multispectral data: the CP-SSM explores cross-modal complementary information, while the SP-SSM strengthens modality-invariant shared features, ultimately achieving optimal feature fusion through the FF-SSM. This design simultaneously focuses on modality-specific complementary features and modality-invariant shared features, establishing a comprehensive cross-modal feature collaboration mechanism that significantly enhances the representational capability of multispectral data.
@@ 3.3.1 CP-SSM Module
Figure 7: The details of the CP-SSM module. The feature maps ($F_{V},F_{T}$) are first reshaped into a sequence ($x_{i}^{V}$, $x_{i}^{T}$) by row and column scanning, and generated $B,C,\Delta$ through a Linear layer. Secondly, we perform a cross-modal complementary features interaction by exchanging the C of the two branches. Finally, the cross-modal complementary features interaction is conducted by the SSM module to generate the output ($\widetilde{F}_{V},\widetilde{F}_{T}$).
@@
3.3.2 FF-SSM Module
Although the CP-SSM module can capture implicit complementary relationships between modalities, it still has limitations at the feature fusion level. To address this, we propose an innovative FF-SSM module that achieves deep feature interaction through bidirectional sequence modeling. Specifically, given the feature maps $F_{1}$ and $F_{2}$ output by CP-SSM, this module employs a bidirectional heterogeneous sequence construction strategy: concatenating the features in different orders ($F_{1},F_{2}$ and $F_{2},F_{1}$). This design not only expands the model’s receptive field but also enhances the diversity of feature representation. In implementation, the features are first unfolded and concatenated for the two directional sequences, followed by generating corresponding state space model parameters ($B_{1},C_{1},\Delta_{1}$ and $B_{2},C_{2},\Delta_{2}$) through linear transformation layers. Finally, by fusing the output features ($F_{12}$ and $F_{21}$) from the bidirectional SSM paths, the module achieves thorough cross-modal feature interaction and adaptive fusion, significantly improving detection accuracy. Unlike the Transformer, which divides the sequence into smaller chunks and overlooks intra-chunk information, the FF-SSM module retains the original sequence information as input. The FF-SSM module effectively preserves the detailed information within both modalities, ensuring richer and more precise feature representation, which is formulated in Equation 7:
$\widehat{F}$ $=\psi_{Merge}\left(F_{12},F_{21}\right)$ (7)
@@ 3.4 Loss Function
where $\mathcal{L}_{\text{QFL}}$, $\mathcal{L}_{\text{L1}}$, $\mathcal{L}_{\text{GIoU}}$, $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{Focal}}$ are the quality focal loss, L1 loss, GIoU loss, cross-entropy loss and focal loss. Hyperparameters $\lambda$s are the weighting factor for each loss term.

### v2

@@ Abstract
Modern multispectral feature fusion for object detection faces two critical limitations: (1) Excessive preference for local complementary features over cross-modal shared semantics adversely affects generalization performance; and (2) The trade-off between the receptive field size and computational complexity present critical bottlenecks for scalable feature modeling. Addressing these issues, a novel Multispectral State-Space Feature Fusion framework, dubbed MS2Fusion, is proposed based on the state space model (SSM), achieving efficient and effective fusion through a dual-path parametric interaction mechanism. More specifically, the first cross-parameter interaction branch inherits the advantage of cross-attention in mining complementary information with cross-modal hidden state decoding in SSM. The second shared-parameter branch explores cross-modal alignment with joint embedding to obtain cross-modal similar semantic features and structures through parameter sharing in SSM. Finally, these two paths are jointly optimized with SSM for fusing multispectral features in a unified framework, allowing our MS2Fusion to enjoy both functional complementarity and shared semantic space. Benefiting from the design of the dual-branch SSM, our approach simultaneously inherits the computational efficiency and the global receptive field, significantly improving the performance of multispectral object detection. Our experiments demonstrate that MS2Fusion consistently outperforms strong baselines: it improves mAP@0.5 by 3.3% for object detection (FLIR), boosts mIoU by 2.8% for RGB-T semantic segmentation (MFNet), and reduces MAE by 0.7% for salient object detection (VT5000). Notably, without requiring task-specific modifications, MS2Fusion achieves new state-of-the-art results across multiple multispectral perception tasks, demonstrating its remarkable generalization capability. The source code is available at https://github.com/61s61min/MS2Fusion.git.
@@ 1 Introduction
Despite recent advances in multispectral object detection, existing methods predominantly focus on learning complementary features across modalities while neglecting the exploration of shared feature representations and inherent structural similarities. Most approaches adopt a single, direct fusion strategy, which often suppresses discriminative yet weaker features and leads to significant information loss. Moreover, while CNN-based methods struggle with limited receptive fields for capturing cross-modal context, Transformer-based techniques suffer from computational inefficiency and performance degradation with long sequences. In speech recognition [10] and image compression [11], the efficiency of feature representation directly determines system performance - a principle equally applicable to multispectral detection. These limitations highlight the need for a more adaptive and efficient fusion paradigm that not only bridges modal discrepancies but also enhances single-modal features through shared representations, ultimately improving robustness in challenging environments.
@@
As demonstrated in Figure 3, our ERF analysis reveals that MS2Fusion achieves superior spatial coverage compared to both CNN and Transformer, successfully integrating local details with global context. This capability addresses the fundamental constraint of standard CNN in modeling long-range dependencies while maintaining computational efficiency. Table 1 also provides our component replacement experiments on the MS2Fusion module in the YoloV5 framework on FLIR dataset(Figure 3), which yield three key findings: (1) Compared to Transformer, MS2Fusion achieves a 66.6% reduction in FLOPs and 70.4% fewer parameters while delivering better detection accuracy (+8.3 mAP50 points); (2) When benchmarked against CNN baseline, it maintains 26.0% lower computational costs and 18.4% parameter reduction, while achieving significant improvements of +9.0 mAP50 points; (3) Across all complexity-accuracy trade-off metrics, MS2Fusion consistently outperforms both baseline models. Notably, our approach achieves processing speed comparable to CNN (with only a 2.1 frames difference) while being significantly faster than Transformer (by 10.5 frames). These systematic experiments provide conclusive evidence that MS2Fusion successfully breaks the traditional efficiency-accuracy trade-off, establishing new state-of-the-art performance in multispectral object detection.
Table 1: Comparison of different multispectral feature fusion methods(YoloV5 framework, FLIR dataset)
GFLOPs Params (M) FPS mAP@0.5 mAP@0.75 mAP
CNN 190.3 159.7 30.8 74.3 23.8 32.5
Transformer 421.9 440.6 19.2 75.0 24.1 33.4
MS2Fusion 140.8 130.3 29.7 83.3 33.0 40.3
@@
    To address the limitations of existing methods that either overlook cross-modal interactions or overly rely on complementary features,
    MS2Fusion introduces a novel state-space framework. Its CP-SSM captures implicit feature complementarity while SP-SSM enhances shared representations, achieving an optimal balance between modality-specific and cross-modal learning.
@@
    Motivated by the need of combining global context modeling with computational efficiency, our FF-SSM mechanism uniquely achieves Transformer-level receptive fields while maintaining CNN-like efficiency. This breakthrough resolves the longstanding trade-off between performance and computational cost in multimodal systems.
@@
    MS2Fusion architecture offers flexible multimodal integration, supporting various detection frameworks and backbones. It seamlessly processes RGB, thermal, and other modalities, enabling broad application across different vision tasks without architectural changes.
@@ 3.1 State Space Model (SSM)
SSM [12] combines the advantages of RNNs and CNNs for efficient handling of long dependencies. It transforms the input sequence $x(t)$ into an intermediate state $h(t)$ through a state equation, then generates the output $y(t)$ via an output equation. SSM is typically represented as a linear ODE:
@@
To adapt SSM for deep learning, it is discretized as:
@@
${\overline{C}}$ $=\textbf{C}$
After discretization [38], SSM uses convolution operations for parallel computation. To overcome the issue of fixed parameters, SSM introduces dynamic adjustments, allowing the model to flexibly handle complex data and improve long-sequence modeling.
Figure 4: Details of SSM, where the red, green and black lines correspond to the equations of Equation 2, respectively. (For an $L\times d$ dimensional input $x_{1},x_{2},...,x_{L}$}.) The input sequence is linearly projected to generate parameters for the state equation, followed by recursive computation via matrices A, B, and C, with an optional skip connection D. The output sequence retains the same dimensionality as the input.
As shown in Figure 4, the SSM processes the input sequence ($\{x_{1},x_{2},...,x_{L}\}\in\mathbb{R}^{L\times d}$) through a structured transformation composed of three key operations. First, the input undergoes a linear projection to generate parameters for the state space equation. This is followed by a recursive state update governed by four fundamental matrices: the current input is initially processed by matrix B, which maps the input to the state space, and then added to the hidden state updated by matrix A, producing a new hidden state. For the output, the new hidden state is remapped to the output via matrix C. Additionally, D provides an optional skip connection that bypasses the state transformation, directly adding the input to the output, which helps mitigate gradient vanishing during training. This recurrent computation generates intermediate outputs $\{y_{1},y_{2},...,y_{L}\}\in\mathbb{R}^{L\times d}$ [38], where each $y_{i}\in\mathbb{R}^{1\times d}$ is derived sequentially.
Figure 5: Overview of the model architecture. It consists of three main stages: (1) feature extraction with two backbone networks; (2) cross-modal feature fusion of $P3$, $P4$ and $P5$ with MS2Fusion module; (3) the detection results are generated through the Neck and Head layers. In our experiments, two distinct detection heads ( CoDetr and YOLOv5) are evaluated independently. The MS2Fusion module employs a dual-branch architecture to process the features of two modalities, $F_{V}$ and $F_{T}$. CP-SSM learns cross-modal complementary features, while SP-SSM extracts shared features. Finally, $F$-SSM performs single brach feature enhancement and cross-modal fusion, outputting the fused feature $F_{fused}$.
@@ 3.2 The Proposed Model Architecture
As shown in the upper part of Figure 5, the proposed framework is structured into three main stages. Initially, the backbone network extracts features independently from RGB and thermal images. Subsequently, cross-modal MS2Fusion feature fusion integrates features from different stages. Finally, object localization and regression are performed using the detection head to derive the final detection outcomes. During the feature fusion stage, we selectively merge features from three distinct levels. P3 layer captures detailed surface information, while P5 encapsulates higher-level semantic feature. By separately fusing shallow and deep features, our approach effectively concentrates on detailed information critical for cross-modal fusion, which is formulated in Equation 3:
@@ 3.3 Multispectral State Space Feature Fusion (MS2Fusion)
As shown in the lower part of Figure 5, the MS2Fusion module consists of three core components: CP-SSM, SP-SSM, and FF-SSM. CP-SSM module achieves global feature integration of RGB and thermal modalities through a dynamic parameter interaction mechanism, enhancing cross-modal contextual awareness while preserving modality-specific characteristics. This module innovatively employs an implicit parameter crossover strategy to effectively mine and reinforce complementary features between the two modalities. CP-SSM allows adaptive parameter modulation based on inter-modal relationships, enabling finer-grained complementary feature extraction. In contrast, the SP-SSM is designed to learn shared, modality-invariant representations that are consistent across both RGB and thermal inputs. All parameters in SP-SSM are constrained to be identical between modalities, forcing the model to discard modality-specific noise and to focus on common semantic information, such as object contours, shapes, and scene structure.
The key distinction between CP-SSM and SP-SSM lies in their parameter handling strategies and functional objectives. The CP-SSM employs dynamic parameter interaction mechanisms where modality-specific parameters are selectively exchanged or combined to capture complementary cross-modal dependencies, while the SP-SSM strictly maintains identical parameters across modalities to enforce feature consistency in shared spaces.
Finally, the FF-SSM modules further integrate these learned features in a hierarchical manner, adaptively combining complementary and shared information to achieve more effective cross-modal fusion.
The innovation of MS2Fusion lies in its systematic utilization of the dual characteristics of multispectral data: the CP-SSM explores cross-modal complementary information, while the SP-SSM strengthens modality-invariant shared features, ultimately achieving optimal feature fusion through the FF-SSM. This dual-path design explicitly decouples the learning of complementary and shared features: CP-SSM’s parameters are optimized for cross-modal differentiation, whereas SP-SSM’s parameters are optimized for cross-modal unification. By simultaneously focusing on modality-specific complementary features and modality-invariant shared features, the framework establishes a comprehensive cross-modal feature collaboration mechanism that significantly enhances the representational capability of multispectral data.
@@ 3.3.1 CP-SSM Module
Figure 6: The details of the CP-SSM module. The feature maps ($F_{V},F_{T}$) are first reshaped into a sequence ($x_{i}^{V}$, $x_{i}^{T}$) by row and column scanning, and generated $B,C,\Delta$ through a Linear layer. Secondly, we perform a cross-modal complementary features interaction by exchanging the C of the two branches. Finally, the cross-modal complementary feature interaction is conducted by the SSM module to generate the output ($\widetilde{F}_{V},\widetilde{F}_{T}$).
@@ 3.3.2 SP-SSM Module
3.3.3 FF-SSM Module
Although the CP-SSM module can capture implicit complementary relationships between modalities, it still has limitations at the feature fusion level. To address Mamba’s inherent feature forgetting while preserving its computational efficiency, we propose a bidirectional architecture combining complementary forward ($F_{1},F_{2}$) and reverse ($F_{2},F_{1}$) processing paths. This innovative design systematically compensates for unidirectional modeling limitations through direction-adaptive parameter generation, where the reverse path specifically aims to capture potential feature loss in the forward pass. Unlike traditional Transformers that suffer from information loss due to chunking operations, our approach not only maintains Mamba’s original sequential dependency and linear complexity advantages but also significantly enhances feature retention through reverse-sequence compensation. The dual-path mechanism enables more comprehensive capture of long-range cross-modal dependencies while dynamically mitigating feature forgetting, achieving robust sequential modeling without sacrificing efficiency. Specifically, given the feature maps $F_{1}$ and $F_{2}$ output by CP-SSM, this module employs a bidirectional heterogeneous sequence construction strategy: concatenating the features in different orders ($F_{1},F_{2}$ and $F_{2},F_{1}$). This design not only expands the model’s receptive field but also enhances the diversity of feature representation. In implementation, the features are first unfolded and concatenated for the two directional sequences, followed by generating corresponding state space model parameters ($B_{1},C_{1},\Delta_{1}$ and $B_{2},C_{2},\Delta_{2}$) through linear transformation layers. Finally, by fusing the output features ($F_{12}$ and $F_{21}$) from the bidirectional SSM paths, the module achieves thorough cross-modal feature interaction and adaptive fusion, significantly improving detection accuracy. The FF-SSM module effectively preserves the detailed information within both modalities, ensuring richer and more precise feature representation, which is formulated in Equation 6:
$\widehat{F}$ $=\psi_{Merge}\left(F_{12},F_{21}\right)$ (6)
@@ 3.4 Loss Function
[…]
@@ 4.1 Dataset and Evaluation Metric
VEDAI [47]: It is designed for multi-class vehicle detection in aerial images, comprising 3,640 objects instances across nine categories, including boats, cars, campers, airplanes, shuttles, tractors, trucks, vans, and others. It contains a total of 1,210 images with a resolution of 1,024 × 1,024, each featuring three RGB color channels and one NIR channel. The dataset is divided into training and test sets in a 9:1 ratio. Annotations, originally in a rotated box format with quadratic coordinates, are converted to a horizontal-box format following the methodology described in [48].